OpenStudy (anonymous):

Complete the square for the binomial: x square+6x

OpenStudy (anonymous):

(x+3)^2

OpenStudy (anonymous):

for f(x)=ax^2+bx, to complete the square take b/2 and then square it. so y=x^2+6x; b=6. 6/2=3^2=9 x^2+6x+9